# IDSelect: A RL-Based Cost-Aware Selection Agent for Video-based Multi-Modal Person Recognition

### v1

@@ 1 Introduction
Figure 1: Top: Traditional multi-modal person recognition methods (e.g., QME [40]) use fixed model combinations for all inputs, while our IDSelect employs an RL-based cost-aware agent to adaptively select complementary models from diverse pools based on input characteristics. Bottom: Accuracy vs. computational cost on CCVID dataset [10]. Our method achieves superior accuracy (95.9%, $1.8\%$) with 92.4% fewer FLOPs than QME.
@@
To address these limitations, we introduce IDSelect, an RL-based cost-aware agent that dynamically selects one pre-trained model per modality for each input. Our key insight is that a learned selection agent can discover complementary, input-conditioned model combinations that outperform static ensembles while utilizing significantly fewer computational resources. Importantly, we choose to select only one model per modality rather than multiple models within the same modality. This design choice is motivated by two key observations: First, models within the same modality tend to learn similar feature representations [16, 25], leading to high redundancy and diminishing marginal returns when combined. Second, true complementarity arises from cross-modal diversity; face, gait, and body capture fundamentally different biometric signatures—whereas selecting multiple face (or gait, or body) models primarily increases computational cost without proportional accuracy gains. By intelligently selecting model combinations based on input characteristics and budget constraints, IDSelect achieves competitive recognition performance at a significantly reduced computational cost.
@@ 2 Related work
Whole-Body Person Recognition. Video-based Multi-modal person recognition integrates complementary biometric traits—face, gait, and body shape—to achieve robust identification in challenging environments [22, 17, 14]. Unlike uni-modal approaches that rely on single biometric sources [5, 18, 7, 10], whole-body systems mitigate modality weaknesses through cross-modal complementarity. For instance, face recognition struggles with extreme pose variations; gait analysis is sensitive to clothing changes, and body shape can be affected by posture—yet their combination provides robustness where individual modalities fail. Recent advances have explored end-to-end pipelines integrating detection, restoration, and recognition [22, 23], demonstrating improved performance in unconstrained settings. However, these systems typically deploy all models simultaneously, leading to substantial computational overhead regardless of input complexity.
@@
Adaptive and Cost-Aware Inference. The computer vision community has developed sophisticated methods for input-dependent computation. Early-exit networks dynamically terminate inference when confidence thresholds are met. Teerapittayanon et al[28] established the foundational BranchyNet framework for confidence-based early exits, while Yang et al[32] demonstrated spatial redundancy exploitation via Resolution Adaptive Networks. Recent work on Adaptive Vision Transformers has pushed boundaries further: Yin et al[34] developed A-ViT, achieving significant throughput improvements with minimal accuracy drop through adaptive token reduction, and Meng et al[24] introduced fine-grained control over computational resources via learned usage policies. In video-based multi-modal settings, Panda et al[26] proposed AdaMML for video understanding, achieving computational reduction through policies that adaptively select modality subsets. However, these adaptive approaches have not been systematically applied to multi-modal biometric recognition.
@@ 3.3 Model Training
where $\sigma(\cdot)$ is the sigmoid function, BCE denotes binary cross-entropy, and $C(\mathbf{a})$ is the average of modality-specific FLOP costs normalized by their modality maxima. The scalar $\lambda$ is a Lagrange multiplier that self-adjusts to satisfy the target budget $C_{\text{target}}$: after each batch, we update $\lambda\leftarrow\max(0,\lambda+\eta(\bar{C}-C_{\text{target}}))$, where $\bar{C}$ is the observed batch-average cost and $\eta$ is a small step size.
@@ 4.1 Experimental Setup
Table 1: Performance comparison on CCVID Configuration 1. IDSelect achieves superior accuracy-efficiency trade-offs compared to single modality baselines, fixed fusion strategies, and quality-guided methods. ’Ada.’ denotes input-adaptive selection. Colored icons in “Comb.” represent modality and model capacity ($\circ$=face, $\square$=body, $\triangle$=gait; fill level increases with model size, from light to heavy). Models marked with an asterisk * are not finetuned on the CCVID dataset, while all other models are finetuned.
@@
Weighted-sum Fusion
Face + Gait Max 693.6 91.4 85.7
Face + Gait Min 11.7 94.4 91.7
Face + Body Max 36.8 93.6 90.0
Face + Body Min 12.8 92.2 86.6
Gait + Body Max 681.8 84.2 78.7
Gait + Body Min 14.1 89.2 86.3
Face + Body + Gait Max 706.1 91.7 88.9
Face + Body + Gait Min 19.4 94.5 93.5
@@ 4.2 Results on CCVID Dataset
Tab. 1 presents results on CCVID Configuration 1, where IDSelect demonstrates exceptional efficiency gains while improving accuracy. IDSelect achieves 95.9% Rank-1 and 94.6% mAP while consuming only 53.7 GFLOPs—a 92.4% reduction compared to QME’s 706.1 GFLOPs with +1.8% higher accuracy. This dramatic computational saving demonstrates that adaptive model selection can discover superior model combinations that heavyweight exhaustive fusion approaches fail to find.
Tab. 2 shows results on CCVID Configuration 2 with a more computationally constrained model pool. IDSelect maintains strong performance with 97.4% Rank-1 and 92.7% mAP at 54.8 GFLOPs, achieving 2.2× computational savings compared to exhaustive Max fusion. Notably, Configuration 2 achieves higher accuracy than Configuration 1 (97.4% vs 95.9% Rank-1) despite operating within tighter computational constraints, demonstrating IDSelect’s adaptability to different model pool characteristics while maintaining consistent efficiency.
@@
Figure 3: Model selection frequency distribution for CCVID configurations. (a) Configuration 1 shows concentrated patterns with IDSelect predominantly selecting adaface_101 + gaitset + ap3d_34 (71.3%) and adaface_101 + gaitbase + ap3d_34 (24.2%). (b) Configuration 2 exhibits more diverse selection patterns, demonstrating adaptability to different model pool characteristics.
Figure 4: Adaptive model selection examples on CCVID. The agent uses lighter models for clear inputs and switches to stronger models for challenging or low-quality cases.
@@ 4.4 Analysis and Discussion
Efficiency Trade-offs. IDSelect consistently discovers optimal efficiency regions that are unavailable to fixed strategies. As shown in Tabs. 1 and 2, while Min”combinations achieve reasonable efficiency but lower accuracy (Face + Body + Gait Min: 94.5–95.1% Rank-1 at 19.4–20.2 GFLOPs), Max” combinations provide varying accuracy at a high computational cost (Tab. 1: Face + Body + Gait Max achieves 91.7% Rank-1 at 706.1 GFLOPs). IDSelect finds superior trade-offs with 95.9–97.4% Rank-1 accuracy at a moderate cost (53.7–54.8 GFLOPs). The method achieves substantial savings: a 13.2× reduction on CCVID Configuration 1 while improving accuracy from 94.1% to 95.9%; a 2.2× reduction on Configuration 2; and a 1.7× savings on MEVID while maintaining competitive performance.

### v2

@@ 1 Introduction
Figure 1: Top: Prior multi-modal person recognition methods (e.g., QME [51]) use fixed model combinations for all inputs, while our IDSelect employs an RL-based cost-aware agent to adaptively select complementary models from diverse pools based on input characteristics. Bottom: Accuracy vs. computational cost on CCVID dataset [11]. Our method achieves superior accuracy (95.9%, $1.8\%$) with 92.4% fewer FLOPs than QME.
@@
To address these limitations, we introduce IDSelect, an RL-based cost-aware agent that dynamically selects one pre-trained model per modality for each input. Our key insight is that a learned selection agent can discover complementary, input-conditioned model combinations that outperform static ensembles while utilizing significantly fewer computational resources. Unlike prior RL-based adaptive inference methods [3, 35] that target single-modality early exit within one network, IDSelect addresses the fundamentally different problem of cross-modal expert selection for retrieval. This introduces unique technical challenges absent in prior work: (1) probe-gallery consistency— the same model must be applied to both probe and gallery sequences per modality for valid similarity computation, since different models produce incompatible embeddings; (2) the combinatorial cross-modal action space spans heterogeneous face, gait, and body pools with distinct failure modes, where we show that cross-modal diversity matters more than intra-modal redundancy. Importantly, we choose to select only one model per modality rather than multiple models within the same modality. This design choice is motivated by two key observations: First, models within the same modality tend to learn similar feature representations [17, 31], leading to high redundancy and diminishing marginal returns when combined. Second, true complementarity arises from cross-modal diversity; face, gait, and body capture fundamentally different biometric signatures—whereas selecting multiple face (or gait, or body) models primarily increases computational cost without proportional accuracy gains. By intelligently selecting model combinations based on input characteristics and budget constraints, IDSelect achieves competitive recognition performance at a significantly reduced computational cost.
@@ 2 Related work
Whole-Body Person Recognition. Video-based multi-modal person recognition integrates complementary biometric traits (face, gait, and body) for robust identification in challenging environments [25, 18, 15]. Unlike uni-modal approaches relying on single biometric sources [6, 19, 8, 11], whole-body systems mitigate modality weaknesses through cross-modal complementarity. Face recognition struggles with extreme pose variations [20, 21, 28]; gait analysis is sensitive to clothing changes [36, 49], and body shape can be affected by posture [37, 32, 29, 27]—yet their combination provides robustness where individual modalities fail. Recent advances have explored end-to-end pipelines integrating detection, restoration, and recognition [25, 26], demonstrating improved performance in unconstrained settings. However, these systems deploy all models simultaneously, leading to substantial computational overhead regardless of input complexity.
@@
Adaptive and Cost-Aware Inference. The computer vision community has developed sophisticated methods for input-dependent computation. Bolukbasi et al. [3] proposed adaptive neural networks that learn policies to route inputs through networks of varying capacity, while Rao et al. [35] applied deep RL to attend to informative frames for video face recognition. Early-exit networks dynamically terminate inference when confidence thresholds are met. Teerapittayanon et al[38] established the foundational BranchyNet framework for confidence-based early exits, while Yang et al[42] demonstrated spatial redundancy exploitation via Resolution Adaptive Networks. Recent work on Adaptive Vision Transformers has pushed boundaries further: Yin et al[44] developed A-ViT, achieving significant throughput improvements with minimal accuracy drop through adaptive token reduction, and Meng et al[30] introduced fine-grained control over computational resources via learned usage policies. In video-based multi-modal settings, Panda et al[33] proposed AdaMML for video understanding, achieving computational reduction through policies that adaptively select modality subsets. However, these adaptive approaches have not been systematically applied to multi-modal biometric recognition.
@@ 3.3 Model Training
where $\sigma(\cdot)$ is the sigmoid function, BCE denotes binary cross-entropy, and $C(\mathbf{a})$ is the average of modality-specific FLOP costs normalized by their modality maxima. We use BCE as a differentiable surrogate for the non-differentiable ranking metrics (Rank-1, mAP): encouraging positive-pair similarities to exceed negative-pair similarities aligns with the ranking objective. The scalar $\lambda$ is a Lagrange multiplier that self-adjusts to satisfy the target budget $C_{\text{target}}$: after each batch, we update $\lambda\leftarrow\max(0,\lambda+\eta(\bar{C}-C_{\text{target}}))$, where $\bar{C}$ is the observed batch-average cost and $\eta$ is a small step size.
@@ 4.1 Experimental Setup
Table 1: Performance comparison on CCVID Configuration 1. IDSelect achieves superior accuracy-efficiency trade-offs compared to single modality baselines, fixed fusion strategies, and quality-guided methods. ’Ada.’ denotes input-adaptive selection. Colored icons in “Comb.” represent modality and model capacity ($\circ$=face, $\square$=body, $\triangle$=gait; fill level increases with model size, from light to heavy). Models marked with an asterisk * are not finetuned on the CCVID dataset, while all other models are finetuned. Heavy/Light denotes the highest/lowest-cost model.
@@
Score-Level Fusion
Face + Gait (Heavy) 693.6 91.4 85.7
Face + Gait (Light) 11.7 94.4 91.7
Face + Body (Heavy) 36.8 93.6 90.0
Face + Body (Light) 12.8 92.2 86.6
Gait + Body (Heavy) 681.8 84.2 78.7
Gait + Body (Light) 14.1 89.2 86.3
Face + Body + Gait (Heavy) 706.1 91.7 88.9
Face + Body + Gait (Light) 19.4 94.5 93.5
@@ 4.2 Results on CCVID Dataset
Tab. 1 presents results on CCVID Configuration 1, where IDSelect demonstrates exceptional efficiency gains while improving accuracy. IDSelect (Adaptive) achieves 95.9% Rank-1 and 94.6% mAP while consuming only 53.7 GFLOPs—a 92.4% reduction compared to QME’s 706.1 GFLOPs with ++1.8% higher accuracy. This dramatic computational saving demonstrates that adaptive model selection can discover superior model combinations that heavyweight exhaustive fusion approaches fail to find. Note that feature-level averaging across modalities is inapplicable in our setting, as different models produce embeddings in incompatible spaces and dimensions; score-level mean fusion (the Light/Heavy rows in Tab. 1) is therefore the appropriate simple baseline. IDSelect (Adaptive) outperforms even the strongest such baseline, F + B + G (Light), which achieves 94.5% Rank-1 at 19.4 GFLOPs. To further distinguish IDSelect from static selection approaches (e.g., NAS), we report IDSelect (Static) in Tab. 1, which deploys only the most frequently selected combination from the learned policy. This fixed combination achieves 94.6%, 1.3% below the adaptive variant’s 95.9%, confirming that per-input adaptive selection provides measurable gains that no fixed combination can match.
[…]
@@
Figure 3: (a) Selection frequency distribution for CCVID configurations. Config. 1 shows concentrated patterns with IDSelect predominantly selecting adaface_101 + gaitset + ap3d_34 (71.3%) and adaface_101 + gaitbase + ap3d_34 (24.2%). Config. 2 exhibits more diverse selection patterns, demonstrating adaptability to different model pool characteristics. (b) Selection probability evolution during training (Config. 2). Body converges to TF-CLIP; Gait learns GaitSet/GaitBase preference; Face maintains input-adaptive diversity.
@@ 4.4 Analysis and Discussion
Selection Policy Evolution. Fig. 3 (b) visualizes how selection probabilities evolve during training on Config. 2. Body selection quickly converges to TF-CLIP (the strongest body model), while Gait develops a clear preference for GaitSet/GaitBase. Face selection remains diverse—the policy selects different face models for different inputs (e.g., AdaFace18 $\sim$40%, AdaFace50 $\sim$53% in Fig. 3), so the marginal probability stays balanced while per-input selection is decisive. This confirms input-adaptive rather than random behavior.
RL Policy Comparison. We chose Actor-Critic for its simplicity and stable integration with the Lagrangian budget controller. To validate this choice, we additionally tested PPO: it achieves comparable results (96.1% Rank-1 vs. 95.9%) but adds clipping complexity unnecessary for our small action space (27 combinations). A3C requires asynchronous infrastructure; GRPO targets LLM alignment with group-relative rewards, a fundamentally different setting. Our framework is modular and the base RL algorithm can be swapped without architectural changes.
[…]
Memory Analysis. In full dynamic mode, the 9-model pool (3 per modality) requires $\sim$5.2 GB total VRAM, loaded once at initialization with no runtime swapping. The selection agent adds <<24 MB overhead, modest relative to the 92.4% FLOP reduction, which provides substantial latency and energy savings. For reference, the ablation IDSelect (Static) in Tab. 1 shows that a single fixed combination recovers most gains (94.6% vs. 95.9% Rank-1) at standard 3-model memory, quantifying the 1.3% cost of forgoing adaptivity.